find the trinomial that represents the area of the square if the measure of the side is (7b-3)

Answer by ewatrrr(24785)

 
Hi,
measure of the side os a square is (7b-3)
A(square)= s*s = (7b-3)(7b-3) = 49b^2 - 42b + 9
F First terms
O Outside terms
I Inside terms
L Last terms